OpenStudy (anonymous):

What is the value of u + v + w when u = –10, v = 4, and w = 8? A. –22 B. 2 C. 14 D. 22

OpenStudy (anonymous):

plzzz I need it now!!! :)

OpenStudy (anonymous):

The answer B

OpenStudy (anonymous):

-10+4+8=2

OpenStudy (anonymous):

really? I think it would be -22

OpenStudy (blackfireskull):

no. hype is right. -10+4= -6 -6+8=2

OpenStudy (anonymous):

-10+4=-6 -6+8=2

OpenStudy (anonymous):

thanks ya'll could ya'll get one more? Which statements are true? Choose all answers that are correct. A. –13 + 25 = 10 + (–22) B. 20 + (–6) > 31 + (–19) C. 8 + (–12) < –17 + 14 D. –11 + (–15) > 6 + (–29)

OpenStudy (anonymous):

its B 20+(-6)=14 14>12 31+(-19)=12

OpenStudy (anonymous):

@@lille_parish